Calculate the mass of solid sodium carbonate when 16.8 grams of sodium hydrogencarbonate is heated strongly until there is no further change. (3pts) 2NaHCO3(s) -- Na2CO3(s) + CO2 (g) +H2O (l) *

1
Expert's answer
2021-11-17T07:26:02-0500

"2NaHCO_3 (heat) \to Na_2CO_3+CO_2+H_2O"

Moles of sodium hydrogen carbonate "=\frac{16.8}{23+1+12+16+16+16}=0.2"

Mole ratio, sodium hydrogen carbonate : sodium carbonate=2:1

Moles of sodium carbonate = "0.2\/2=0.1"

Mass of sodium carbonate = no. of mole "\times" molar mass="0.1\times (23+23+12+16+16+16)=10.6g"